OpenStudy (anonymous):

Factor completely p4^4 – 8p^2 + 16

OpenStudy (anonymous):

Is that 4p^4?

OpenStudy (anonymous):

Yes I mean, no, I'm sorry

OpenStudy (anonymous):

It's p^4 **

OpenStudy (anonymous):

(2p^2)^2 - 2(2p^2)(4) + (4)^2 = (2p^2 - 4)^2

OpenStudy (anonymous):

p^4-8p^2+16 (p^2-4)^2

OpenStudy (anonymous):

Then the answer is (p^2 - 4)^2 = (p^2 - 2^2) = (p + 2)(p - 2)

OpenStudy (anonymous):

sorry, small mistake, ((p + 2)(p - 2))^2

OpenStudy (anonymous):

p^2 - 4 can be further factorised yellowranger.

OpenStudy (anonymous):

Oh, good call. Thanks for noticing.